# Semantic Tree Inference on Text Corpa using a Nested Density Approach together with Large Language Model Embeddings

### v1

@@ 3.5 Label Transcription
The theses-fr-en dataset is annotated and contains a research field (discipline, in French) for each abstract. However, these fields can in many cases be freely chosen by the thesis author, resulting in 13,599 unique discipline labels. In a first test we cross annotated these unique discipline labels into a binary classification of either humanities or natural sciences. We then proceeded to obtain a more coherent and standardized annotation, we used the gpt-oss-20b model to classify all 13,599 labels according to the OECD Fields of Science (FOS) 2002 classification [oecd-fos]. The resulting transcribed labels were then used to color the trees inferred from this dataset, allowing us to assess whether specific scientific fields tend to separate into distinct and meaningful branches.
@@ 3.6 Tree Coloration and Visualization
Tree coloration was performed using utilities provided by MNHN-Tree-Tools [mnhn-tree-tools]. In particular, we employed the tree_map_for_split_set tool to map predefined clusters or labeled subsets onto the corresponding nodes of the hierarchical tree. The resulting color maps were generated for visualization using the newick-utilities package [newick-utilities], a software suite for the manipulation and visualization of phylogenetic trees.
@@ 4.1 Efficiency of PCA and the $L_{2}$-norm Compared to Cosine Distance
In contrast, applying Principal Component Analysis (PCA) prior to clustering dramatically improved the quality and expressiveness of the resulting trees. Experiments with 10, 5, and 2 principal components showed that lower-dimensional PCA representations consistently produced more “spread out” trees with clearer branching and more detectable fine-scale semantic substructure. We suggest that this effect is due to the inherent ability of PCA to yield principal components that maximize variance and that this feature selection process has a positive effect on our clustering algorithm. A similar effect was noted previously in the algorithmic supplement of [mnhn-tree-tools].
A comparison between trees obtained with cosine distance and with the $L_{2}$-norm applied to PCA-reduced embeddings is shown in Figure 3. From left to right, the trees become more hierarchically detailed and reveal progressively finer semantic subclusters. For these illustrative examples, the parameters were chosen using a rudimentary bisection search to produce visually optimized trees. The parameter settings were as follows: for cosine distance, initial $\epsilon=0.275$, $\Delta\epsilon=0.001$, and $\mathrm{minpts}=5$; for the $L_{2}$-norm on 10 principal components, initial $\epsilon=0.091$, $\Delta\epsilon=0.0001$, $\mathrm{minpts}=5$; and for the $L_{2}$-norm on 2 principal components, initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$. The trees shown in 3 were built from embeddings obtained form the Qwen3-Embedding-8B model.
Figure 3: A comparison of trees built from the 20 Newsgroups dataset under different distance measures
@@ 4.2 Detection of Topics and Semantics
Depending on the dataset, we applied different evaluation strategies to assess whether, and to what extent, the method proposed in this work produces meaningful and reliable results, given the computational constraints of our setup. The following subsections outline these evaluations and illustrate the behavior of our hierarchical density-based tree construction across diverse text corpora.
@@
To evaluate whether our tree construction method recovers meaningful semantic structure, we applied it to the 20 Newsgroups dataset. The resulting trees were colored according to posting labels, as shown in Figure 4. For each newsgroup we generated a separate color map: posts from the target newsgroup are shown in blue, and posts from all other newsgroups in orange. Both color intensities follow a logarithmic scale proportional to the number of posts in each cluster, with white indicating the absence of posts of a given type.
The inferred tree, as highlighted in figure 4, recovers many interesting relationships between the newsgroups:
@@
  • Computer-related groups. The comp.* groups, including comp.sys.mac.hardware, comp.windows.x, comp.graphics, and comp.os.ms-windows, form a coherent subtree corresponding to computer hardware and software discussions. Some overlap with sci.electronics is visible, consistent with semantically similar content.
@@
Overall the structure seems to intuitively make sense, outlining intervoven fields of discussion. Importantly, our toolchain does not merely cluster newsgroups into isolated groups; it also conveys the relative proximity and semantic distance between them. The resulting tree therefore serves as a compact visual instrument that captures both fine-grained cluster structure and the broader topology of topic similarity within the dataset, all within a single interpretable representation.
Figure 4: Representations of the tree obtained from the 20 Newsgroups dataset. Trees were built from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] embeddings. Parameters used to build tree were: initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$.
@@ 4.2.2 IMDB 50K Reviews
The IMDB 50K Reviews dataset was used to evaluate whether our tree construction method, together with the underlying embedding models, is able to separate reviews by emotional polarity. We therefore built a tree from this dataset following the same procedure as for the other corpora. The resulting tree was then colored analogously to the 20 Newsgroups experiments: clusters containing predominantly positive reviews were assigned a blue shading, while clusters containing predominantly negative reviews were shaded in orange. As before, color intensity follows a logarithmic scale, with white indicating the absence of reviews of a given label and deeper blue or orange indicating higher concentration. Overlaps appear in darker, greenish to black tones.
The resulting tree is shown in Figure 5. Compared to the other datasets, the IMDB tree appears noticeably less structured, suggesting that the underlying corpus is more homogeneous and lacks strong semantic subdivisions beyond sentiment. Nonetheless, we observe a few small subtrees near the periphery of the tree in which positive reviews form a clear majority. Additional localized regions with a majority of positive or negative reviews can also be identified. We speculate that these structures arise from repetitive linguistic patterns that occur more frequently in either positive or negative reviews. However, they do not form large coherent branches, and overall the dataset does not separate cleanly by sentiment.
This observation suggests that the embedding models used: Qwen3-Embedding-8B [qwen3embedding] and SFR-Embedding-Mistral [SFRembedding] encode emotional polarity only to a limited extent. The results highlight the importance of choosing embedding models appropriate for the expected semantic structure when constructing trees.
We also attempted to annotate the IMDB tree with the Llama-4-Scout-17B-16E-Instruct model [meta2024llama4]. The goal was to obtain genre-level descriptions of tree regions, under the assumption that certain movie genres might correlate with predominantly positive or negative reviews. This attempt was unsuccessful: the model did not produce reliable or meaningful genre assignments, and more fundamentally, such information appears to be largely absent from the review text itself. As such, deriving movie genres from this dataset seems infeasible with current methods.
Figure 5: The IMDB 50K Reviews dataset colored by sentiments: The tree was built using the first two principal components derived from SFR-Embedding-Mistral [SFRembedding] embeddings. Parameters yield to the tree building algorithm were: initial $\epsilon=0.5$, $\Delta\epsilon=0.0001$, $\mathrm{minpts}=5$.
@@ 4.2.3 AG News
Figure 6: The tree obtained from the AG News dataset in different colorations. The tree shown was inferred from the first two principal components derived from SFR-Embedding-Mistral [SFRembedding] embeddings. Parameters yield to the tree building algorithm were: initial $\epsilon=0.225$, $\Delta\epsilon=0.0005$, and $\mathrm{minpts}=5$
We observe that the world and sports categories form clearly separable structures: both branch off early from the root of the tree, resulting in two well-defined subtrees with only minimal contamination from other labels. Around the root, however, a small number of scattered articles appear, suggesting that certain headlines are not easily classifiable by the embedding models. These headlines seem to be very distant from the other news items, only joining the root at a very diffuse density.
@@ 4.2.4 The French Theses Dataset: Theses-fr-en
To further evaluate whether our method can identify and organize broad and fine-grained disciplinary structure, we applied our tree-building pipeline to the Theses-fr-en dataset, which consists of English-language abstracts of French doctoral theses. The dataset contains author-supplied disciplinary annotations, but these labels are highly heterogeneous and unstandardized, yielding 13,599 unique descriptors. Therefore, as outlined in Section 3.5, we first performed a cross-annotation step to obtain more tangible labels.
As an initial experiment, we mapped the original annotations to a binary classification: humanities versus natural sciences. The resulting tree clearly separated humanities theses from the remainder of the corpus, forming a large, coherent branch as shown in Figure 7A. Encouraged by this result, we proceeded to apply a more detailed transcription based on the OECD Fields of Science and Technology (FOS) 2002 classification scheme [oecd-fos].
In the tree three major branches are easily identifyable. The branch identified as “humanities” in the binary test contains predominantly Social Sciences (5) and Humanities (6) according to the FOS taxonomy. Social sciences, however, extend beyond this primary branch, with small isolated scattered clusters appearing throughout the tree.
@@
Figure 7: Tree generated from the French Theses theses-fr-en dataset using the first two principal components obtained from the embeddings yield by the SFR-Embedding-Mistral [SFRembedding] model:
@@ 4.2.5 Annotated Institutional Trees: TU Wien and AUB
Our analysis of the two institutional datasets (TU Wien and AUB) differs from the benchmark datasets presented earlier. In the benchmark cases, we projected known labels onto the inferred trees to assess how well embeddings recover an existing classification. In contrast, the institutional datasets are used to examine research structures ab initio: rather than testing whether a predefined taxonomy is recovered, we investigate how the tree itself organizes the institutional research landscape. Although the embedding models were taken trained on external corpora, and not further fine tuned, the resulting tree structure is independent of any curated ontology. This allows us to explore institution-specific disciplinary profiles, structural proximities between fields, and potential emerging clusters or cross-disciplinary connections. The results demonstrate that our method can capture meaningful organizational patterns without relying on predefined categories.
  • At the American University of Beirut, medicine and health-related disciplines dominate the research corpus, reflecting the strong presence of the university hospital. This dominance is visible in the tree structure, where the entire upper half of the circular tree forms the largest and most cohesive subtree. Within this region, several well-defined subbranches correspond to specific healthcare domains, including health and nutrition, health policy and systems research, and public health and epidemiology.
    A distinct humanities subtree is also present, characterized by numerous subbranches labeled as Middle Eastern Studies and Conflict Resolution. In addition, an engineering subtree that covers engineering, physics, and mathematics, is clearly visible. Notably, this engineering subtree merges with the humanities subtree at a higher density than with the medical subtree, suggesting a closer structural relationship between engineering and humanities research topics within this corpus than might be expected, which might be characteristic of the university’s liberal arts tradition.
@@
Figure 8: Tree generated from the AUB dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] model embeddings. The tree was built with the following parameters: initial $\epsilon=0.0045$, $\Delta\epsilon=0.00001$ and $\mathrm{minpts}=5$.
Figure 9: Tree generated from the TU Wien dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] model embeddings. The tree was built with the following parameters: initial $\epsilon=0.0028$, $\Delta\epsilon=0.000001$ and $\mathrm{minpts}=5$.
@@ 5.1 Insights from Benchmark Datasets
The AG news dataset similarly shows that World and Sports articles form well separated subtrees, but that Business and Science/Technology headlines exhibit a substantial overlap. This result can intuitively explained by the topical similarities between these two categories. Our tree based method as such creates an intrinsic semantic organization derived from LLM-embeddings rather then enforcing a predefined label boundary.
The French doctoral theses dataset provided a further test for our algorithm. This dataset included “discipline” annotations. However these are rather arbitrarily chosen and unstandarized. We transcribed these labels first into a binary annotation of either being part of natural sciences or humanities. To our surprise the theses abstracts of labeled humanities formed a clear subtree in our tree. Encouraged by these findings we progressed to transcribe the “discipline” labels into the OECD FOS annotation. We colored the tree according to it and found thanks to the hierarchical structure that engineering and medical sciences merge earlier than either social sciences or humanities. A relationship that as such cannot be derived from the FOS classification. At the same time we see that finer grained subtrees do correspond cleanly to the FOS subclasses, indicating that our inferred structure captures conceptual relationships that differ from established disciplinary boundaries.
@@
The IMDB 50K Reviews dataset yields a very homogeneous tree, reflecting that the embedding models used in this study are not able to distinguish the semantic differences between the movie reviews in a meaningful way. The labels of this dataset are largely dispersed across the inferred tree, outlining the limitations of our approach, where the embeddings as outlined here do not encode as here sentiment or an other required topic separator.

### v2

@@ 3.6 Label Transcription
The theses-fr-en dataset is annotated and contains a research field (discipline, in French) for each abstract. However, these fields can in many cases be freely chosen by the thesis author, resulting in 13,599 unique discipline labels. In a first test we cross annotated these unique discipline labels into a binary classification of either humanities or natural sciences. We then proceeded to obtain a more coherent and standardized annotation, we used the gpt-oss-20b model to classify all 13,599 labels according to the OECD Fields of Science (FOS) 2002 classification [27]. The resulting transcribed labels were then used to color the trees inferred from this dataset, allowing us to assess whether specific scientific fields tend to separate into distinct and meaningful branches.
@@ 3.7 Tree Coloration and Visualization
Tree coloration was performed using utilities provided by MNHN-Tree-Tools [13]. In particular, we employed the tree_map_for_split_set tool to map predefined clusters or labeled subsets onto the corresponding nodes of the hierarchical tree. The resulting color maps were generated for visualization using the newick-utilities package [16], a software suite for the manipulation and visualization of phylogenetic trees.
@@ 4.1 Efficiency of PCA and the $L_{2}$-norm Compared to Cosine Distance
In contrast, applying Principal Component Analysis (PCA) prior to clustering dramatically improved the quality and expressiveness of the resulting trees. Experiments with 10, 5, and 2 principal components showed that lower-dimensional PCA representations consistently produced more “spread out” trees with clearer branching and more detectable fine-scale semantic substructure. We suggest that this effect is due to the inherent ability of PCA to yield principal components that maximize variance and that this feature selection process has a positive effect on our clustering algorithm. A similar effect was noted previously in the algorithmic supplement of [13].
A comparison between trees obtained with cosine distance and with the $L_{2}$-norm applied to PCA-reduced embeddings is shown in Figure 3. From left to right, the trees become more hierarchically detailed and reveal progressively finer semantic subclusters. For these illustrative examples, the parameters were chosen using a rudimentary bisection search to produce visually optimized trees. The parameter settings were as follows: for cosine distance, initial $\epsilon=0.275$, $\Delta\epsilon=0.001$, and $\mathrm{minpts}=5$; for the $L_{2}$-norm on 10 principal components, initial $\epsilon=0.091$, $\Delta\epsilon=0.0001$, $\mathrm{minpts}=5$; and for the $L_{2}$-norm on 2 principal components, initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$. The trees shown in 3 were built from embeddings obtained from the Qwen3-Embedding-8B model.
Figure 3: A comparison of trees built from the 20 Newsgroups dataset under different distance measures.
@@ 4.2 Detection of Topics and Semantics
To evaluate whether the proposed hierarchical density-based method recovers meaningful semantic structure, we combine qualitative analysis of inferred trees with quantitative evaluation against available ground truth labels. Since the output of our method is a hierarchy rather than a single flat partition, evaluation must account for the fact that clusters emerge, merge, and disappear across tree layers.
For datasets with reference labels, we therefore project the hierarchy onto a sequence of flat clusterings, one per tree layer, and assess their agreement with the ground truth using the Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI). ARI measures the agreement between two partitions while correcting for chance, whereas NMI quantifies the amount of shared information between cluster assignments and labels, normalized to allow comparison across different numbers of clusters. Both measures are well established in information retrieval and clustering evaluation and are complementary: ARI is sensitive to exact cluster membership, while NMI captures coarser label alignment even when cluster boundaries are imperfect.
Across the benchmark datasets, the evolution of ARI and NMI as a function of tree depth reveals dataset-specific behavior that reflects differences in semantic structure rather than methodological artifacts. For the 20 Newsgroups and AG News corpora, agreement with ground-truth labels increases from the finest partitions toward intermediate density levels, reaches a maximum, and then exhibits sharp decreases followed by extended plateaus as clusters merge at coarser scales. This pattern indicates that semantic alignment with predefined categories is strongest at intermediate resolutions, where fine-grained lexical variation has been absorbed while broader thematic distinctions are still preserved. Details are outlined in Figure 7.
In contrast, the IMDB 50K Reviews dataset exhibits a different trend. Here, the highest ARI and NMI values occur closer to the leaves of the tree, within relatively fine-grained branches rather than near the root as shown in Figure 7. As density constraints are relaxed and clusters merge, agreement with sentiment labels decreases steadily, and no dominant peak is observed at intermediate or coarse levels. This behavior reflects the largely homogeneous semantic structure of the corpus, where sentiment is encoded as a subtle, distributed signal rather than as a dominant organizing principle. Consequently, sentiment-aligned clusters emerge only locally and do not persist as stable large-scale structures.
Taken together, these results demonstrate that the proposed hierarchical density-based trees adapt to the intrinsic semantic organization of each dataset. Rather than enforcing a fixed optimal resolution, the method exposes multiple meaningful semantic scales, allowing the analyst to identify the levels at which different types of structure-topical, thematic, or affective, are most pronounced.
Importantly, at high-density (outer) layers, only compact and highly coherent clusters are present, while diffuse documents remain unassigned. As density constraints are relaxed, clusters merge and eventually encompass the full dataset. ARI and NMI are therefore evaluated as a function of tree depth, allowing us to quantify how semantic agreement evolves across scales rather than at a single, arbitrarily chosen resolution.
@@
We first apply our method to the 20 Newsgroups dataset, which provides a well-studied benchmark with twenty topical labels. Figure 4 shows the inferred tree, colored according to individual newsgroup labels.
Quantitatively, ARI and NMI peak at intermediate tree depths, indicating that the strongest alignment between inferred clusters and newsgroup labels occurs neither at the finest nor at the coarsest resolution. At very high densities, clusters are too fragmented to recover complete topics, while at low densities, semantically distinct groups merge into broader themes. This behavior is expected for a hierarchical method and confirms that meaningful topic structure emerges at specific semantic scales.
Qualitatively, the tree recovers intuitive relationships between newsgroups:
@@
  • Computer-related groups. The comp.* groups, including
    comp.sys.mac.hardware, comp.windows.x,
    comp.graphics, and
    comp.os.ms-windows, form a coherent subtree corresponding to computer hardware and software discussions. Some overlap with sci.electronics is visible, consistent with semantically similar content.
@@
These qualitative observations align with the quantitative ARI/NMI profiles: tree regions that correspond to stable semantic groupings also contribute most strongly to the metric peaks. Crucially, the tree representation conveys not only cluster membership but also relative semantic proximity between topics, which is not captured by flat clustering alone.
Figure 4: Representations of the tree obtained from the 20 Newsgroups dataset. Trees were built from the first two principal components obtained from Qwen3-Embedding-8B [34] embeddings. Parameters used to build the tree were: initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$.
@@ 4.2.2 IMDB 50K Reviews
We next evaluate the IMDB 50K Reviews dataset, which provides binary sentiment labels (positive vs. negative). Unlike topical datasets, this corpus is semantically homogeneous, and sentiment is expressed largely through subtle linguistic cues.
ARI and NMI values remain comparatively low across all tree depths, with only a modest peaks at layers still in the high dense leaf node range. This indicates that the embedding models used encode sentiment only weakly relative to other semantic dimensions. The inferred tree, shown in Figure 5, reflects this quantitatively observed behavior: the structure is shallow, with few large subtrees and only localized regions dominated by either positive or negative reviews. As before, color intensity follows a logarithmic scale, with white indicating the absence of reviews of a given label and deeper blue or orange indicating higher concentration. Overlaps appear in darker, greenish to black tones.
Small terminal branches enriched in positive or negative sentiment do exist, suggesting that repetitive phrasing or stylistic patterns occasionally induce local clustering. However, no large-scale sentiment-based separation emerges. These findings highlight an important methodological point: hierarchical semantic trees faithfully reflect the information encoded in the embedding space and should not be expected to recover distinctions that are not strongly represented by the underlying model.
This observation suggests that the embedding models used: Qwen3-Embedding-8B [34] and SFR-Embedding-Mistral [22] encode emotional polarity only to a limited extent. The results highlight the importance of choosing embedding models appropriate for the expected semantic structure when constructing trees.
We also attempted to annotate the IMDB tree with the Llama-4-Scout-17B-16E-Instruct model [1]. The goal was to obtain genre-level descriptions of tree regions, under the assumption that certain movie genres might correlate with predominantly positive or negative reviews. This attempt was unsuccessful: the model did not produce reliable or meaningful genre assignments, and more fundamentally, such information appears to be largely absent from the review text itself. As such, deriving movie genres from this dataset seems infeasible with current methods.
Figure 5: The IMDB 50K Reviews dataset colored by sentiments: The tree was built using the first two principal components derived from SFR-Embedding-Mistral [22] embeddings. Parameters yield to the tree building algorithm were: initial $\epsilon=0.5$, $\Delta\epsilon=0.0001$, $\mathrm{minpts}=5$.
@@ 4.2.3 AG News
Figure 6: The tree obtained from the AG News dataset in different colorations. The tree shown was inferred from the first two principal components derived from SFR-Embedding-Mistral [22] embeddings. Parameters yield to the tree building algorithm were: initial $\epsilon=0.225$, $\Delta\epsilon=0.0005$, and $\mathrm{minpts}=5$
We observe that the world and sports categories form clearly separable structures: both branch off early from the root of the tree, resulting in two well-defined subtrees with only minimal contamination from other labels. This is also reflected by stronger ARI and NMI value peaks than in the other datasets as outlined in Figure 7. Around the root, however, a small number of scattered articles appear, suggesting that certain headlines are not easily classifiable by the embedding models. These headlines seem to be very distant from the other news items, only joining the root at a very diffuse density.
@@
Figure 7: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI) as a function of tree depth for the 20 Newsgroups, IMDB 50K Reviews and AG News dataset. 20 Newsgroups and AG News peak at intermittent clusters indicating that semantic agreement with ground-truth labels is maximized at intermediate density levels rather than at the finest or coarsest partitions.
@@ 4.2.4 The French Theses Dataset: Theses-fr-en
To further evaluate whether our method can identify and organize broad and fine-grained disciplinary structure, we applied our tree-building pipeline to the Theses-fr-en dataset, which consists of English-language abstracts of French doctoral theses. The dataset contains author-supplied disciplinary annotations, but these labels are highly heterogeneous and unstandardized, yielding 13,599 unique descriptors. Therefore, as outlined in Section 3.6, we first performed a cross-annotation step to obtain more tangible labels. As these labels were generated by us we did not apply ARI and NMI evaluations on this tree.
As an initial experiment, we mapped the original annotations to a binary classification: humanities versus natural sciences. The resulting tree clearly separated humanities theses from the remainder of the corpus, forming a large, coherent branch as shown in Figure 8A. Encouraged by this result, we proceeded to apply a more detailed transcription based on the OECD Fields of Science and Technology (FOS) 2002 classification scheme [27].
In the tree three major branches are easily identifiable. The branch identified as “humanities” in the binary test contains predominantly Social Sciences (5) and Humanities (6) according to the FOS taxonomy. Social sciences, however, extend beyond this primary branch, with small isolated scattered clusters appearing throughout the tree.
@@
Figure 8: Tree generated from the French Theses theses-fr-en dataset using the first two principal components obtained from the embeddings yield by the SFR-Embedding-Mistral [22] model:
@@ 4.2.5 Annotated Institutional Trees: TU Wien and AUB
Our analysis of the two institutional datasets (TU Wien and AUB) differs from the benchmark datasets presented earlier. In the benchmark cases, we projected known labels onto the inferred trees to assess how well embeddings recover an existing classification. In contrast, the institutional datasets are used to examine research structures ab initio: rather than testing whether a predefined taxonomy is recovered, we investigate how the tree itself organizes the institutional research landscape. Consequently, ARI and NMI are not applicable. Although the embedding models were trained, by their creators, on external corpora, and not further fine tuned, the resulting tree structure is independent of any curated ontology. This allows us to explore institution-specific disciplinary profiles, structural proximities between fields, and potential emerging clusters or cross-disciplinary connections. The results demonstrate that our method can capture meaningful organizational patterns without relying on predefined categories.
  • American University of Beirut (AUB)
    At AUB, medicine and health-related disciplines dominate the research corpus, reflecting the strong presence of the university hospital. This dominance is visible in the tree structure, where the entire upper half of the circular tree forms the largest and most cohesive subtree. Within this region, several well-defined subbranches correspond to specific healthcare domains, including health and nutrition, health policy and systems research, and public health and epidemiology.
    A distinct humanities subtree is also present, characterized by numerous subbranches labeled as Middle Eastern Studies and Conflict Resolution. In addition, an engineering subtree that covers engineering, physics, and mathematics is clearly visible. Notably, this engineering subtree merges with the humanities subtree at a higher density than with the medical subtree, suggesting a closer structural relationship between engineering and humanities research topics within this corpus than might be expected, which might be characteristic of the university’s liberal arts tradition.
@@
Figure 9: Tree generated from the AUB dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [34] model embeddings. The tree was built with the following parameters: initial $\epsilon=0.0045$, $\Delta\epsilon=0.00001$ and $\mathrm{minpts}=5$.
Figure 10: Tree generated from the TU Wien dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [34] model embeddings. The tree was built with the following parameters: initial $\epsilon=0.0028$, $\Delta\epsilon=0.000001$ and $\mathrm{minpts}=5$.
@@ 5.1 Insights from Benchmark Datasets
The AG news dataset similarly shows that World and Sports articles form well separated subtrees, but that Business and Science/Technology headlines exhibit a substantial overlap. This result can intuitively be explained by the topical similarities between these two categories. Our tree based method as such creates an intrinsic semantic organization derived from LLM-embeddings rather than enforcing a predefined label boundary.
The French doctoral theses dataset provided a further test for our algorithm. This dataset included “discipline” annotations. However these are rather arbitrarily chosen and unstandardized. We transcribed these labels first into a binary annotation of either being part of natural sciences or humanities. To our surprise the theses abstracts of labeled humanities formed a clear subtree in our tree. Encouraged by these findings we progressed to transcribe the “discipline” labels into the OECD FOS annotation. We colored the tree according to it and found thanks to the hierarchical structure that engineering and medical sciences merge earlier than either social sciences or humanities. A relationship that as such cannot be derived from the FOS classification. At the same time we see that finer grained subtrees do correspond cleanly to the FOS subclasses, indicating that our inferred structure captures conceptual relationships that differ from established disciplinary boundaries.
@@
The IMDB 50K Reviews dataset yields a very homogeneous tree, reflecting that the embedding models used in this study are not able to distinguish the semantic differences between the movie reviews in a meaningful way. The labels of this dataset are largely dispersed across the inferred tree, outlining the limitations of our approach, where the embeddings as outlined here do not encode as here sentiment or any other required topic separator.